OpenStudy (anonymous):

The perimeter of a rectangle is 78 cm. The length is 3 greater than twice its width. Find the dimensions.

OpenStudy (deepolisnoob):

l=3+2w p=2l+2w p=2(3+2w)+2w P=6+4w+2w 78=6+6w 72=6w 12=w l=3+2w, l=3+24,l=27 l=27, w=12 27+27+12+12=78

OpenStudy (anonymous):

so the answer is 27 and 12?